# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ 3.3 Training-Free Bayesianization (TFB)
The previous section presents a straightforward Bayesianization scheme for cases where the variational distribution of the full-weight matrix ${\bm{W}}_{0}+\Delta{\bm{W}}$ is constrained to the isotropic Gaussian family, assuming a predetermined value of $\sigma_{q}$. In this section, we describe a practical method for determining $\sigma_{q}$.
@@
A General Bayesianization Framework. Consider an in-distribution “anchor” dataset ${\mathcal{D}}$, an associated evaluation metric $l$, and a permissible performance degradation threshold $\epsilon$. Training-Free Bayesianization (TFB) formulates this as a constrained optimization problem:
@@
where $l({\mathcal{D}}|{\bm{B}}^{\prime},{\bm{M}},{\bm{\Omega}}(\sigma_{q}))=\mathbb{E}_{{\bm{E}}\sim{\mathcal{N}}(0,{\bm{\Omega}})}[l({\mathcal{D}}|{\bm{B}}^{\prime},{\bm{M}}+{\bm{E}})]$ represents the post-Bayesianization performance. This optimization seeks to maximize the noise scale $\sigma_{q}$ applied to model parameters ${\bm{M}}$ while ensuring that the resulting performance degradation (or loss increase) remains within the acceptable threshold $\epsilon$.
@@
Anchor Dataset ${\mathcal{D}}$ and Evaluation Metric $l$. The framework of TFB accommodates various choices of anchor dataset ${\mathcal{D}}$ and evaluation metric $l$ based on practical requirements. We consider two key scenarios (with $N$ being slightly overloaded in its notation):
@@
For supervised dataset ${\mathcal{D}}=\{{\bm{x}}_{n},y_{n}\}_{n=1}^{N}$: The Negative-Log Likelihood (NLL) serves as a natural evaluation metric: $l_{\text{nll}}({\mathcal{D}})=-\tfrac{1}{N}\sum_{n=1}^{N}\log P_{\bm{\theta}}(y_{n}|{\bm{x}}_{n})$, as it directly corresponds to minimizing the variational free energy, as we demonstrate in Sec. 4. The anchor dataset ${\mathcal{D}}$ can be either the original training set used for the LoRA model or an independent calibration dataset, as commonly employed in calibration-based methods (Guo et al., 2017; Zhao et al., 2021). Alternative evaluation metrics such as accuracy or F1 score are also readily applicable. In our experimental setup, to ensure fair comparisons across uncertainty estimation baselines, we use the original training data as ${\mathcal{D}}$ (maintaining the same information access as baselines) and employ accuracy as the evaluation metric.
@@
For unsupervised dataset ${\mathcal{D}}=\{{\bm{x}}_{n}\}_{n=1}^{N}$: Our framework offers substantial flexibility compared to pure calibration methods. One approach is to generate pseudo-labels $\widehat{y}$ using the model before Bayesianization, effectively converting the problem to the supervised case with ${\mathcal{D}}=\{{\bm{x}}_{n},\widehat{y}_{n}\}_{n=1}^{N}$. Additionally, the framework supports alternative evaluation metrics and statistical measures specifically designed for unsupervised data.
Performance Change Tolerance $\epsilon$. The selection of performance change tolerance $\epsilon$ is critical in TFB. While our experiments demonstrate that a fixed $\epsilon=1\%$ for the accuracy metric achieves effective uncertainty estimation across various datasets and LoRA checkpoints (Table 8), an adaptive $\epsilon$ can further enhance the performance of TFB (Table 1).
Many factors can affect the setting of $\epsilon$, among which the most important is the property of the given LoRA checkpoint. For instance, an overfitted LoRA can typically accommodate a larger tolerance $\epsilon$ when using the training dataset (or its subset) as the anchor dataset. Additional characteristics of the data, model, and adaptation tasks can also inform the choice of $\epsilon$. For simplicity and reproducibility in this work, we consider only a binary indicator of LoRA overfitting (which can be readily assessed through ECE when a small validation set is available). We set $\epsilon=1\%$ for overfitted LoRAs and $\epsilon=0.5\%$ for others as the maximum acceptable accuracy degradation.
@@ 3.4 TFB: Final Algorithm
Our final algorithm, presented in Algorithm 1, employs a search strategy to determine the optimal $\sigma_{q}^{*}$ within an initial range $[{\sigma_{q}}_{\text{min}},{\sigma_{q}}_{\text{max}}]$. While binary search would guarantee the optimal solution, we opt for a more efficient parallel approach in practice: we conduct a grid search across 8 different $\sigma_{q}$ values in parallel, construct an approximate function $\widehat{\sigma}_{q}(p)$ through piecewise linear interpolation of the observed performance, and estimate $\sigma_{q}^{*}\approx\widehat{\sigma}_{q}(p_{0}-\epsilon)$, where $p_{0}$ denotes the model’s performance before TFB.
After identifying the optimal $\sigma_{q}^{*}$, we Bayesianize all LoRA layers using this value. For prediction, we sample multiple times from the weight posterior and average the output:
$P_{{\bm{\theta}}}(y|{\bm{x}})$ $=\mathbb{E}_{q({\bm{W}}|{\bm{\theta}})}[P(y|{\bm{x}},{\bm{W}})]$ (8)
$\approx\tfrac{1}{N}\sum_{n=1}^{N}P({\bm{y}}|{\bm{x}},{\bm{W}}_{n}),\quad{\bm{W}}_{n}\sim q({\bm{W}}|{\bm{\theta}}),$
where $q({\bm{W}}|{\bm{\theta}})$ denotes the variational distribution defined in Eqn. 3, and we set the number of test-time samples to $N=10$, following BLoB’s protocol (Wang et al., 2024d).
Algorithm 1 Training-Free Bayesianization (TFB)
0:  ${\mathcal{D}}$: Anchor Dataset;
0:  $\{{\bm{B}},{\bm{A}}\}$: Low-Rank Component;
0:  $l$: Model Evaluation Metric;
0:  $\epsilon$: Performance Change Tolerance;
0:  $[{\sigma_{q}}_{\text{min}},{\sigma_{q}}_{\text{max}}]$: search range of the posterior STD.
1:  Evaluate the original performance: $p_{0}\leftarrow l({\mathcal{D}}|{\bm{B}},{\bm{A}})$.
[…]
4:  while $\sigma_{q}$ not converged do
5:     $\sigma_{q}\leftarrow\nicefrac{{({\sigma_{q}}_{\text{max}}+{\sigma_{q}}_{\text{min}})}}{{2}}$.
[…]
7:     Evaluate the performance:  $p\leftarrow l({\mathcal{D}}|{\bm{B}}^{\prime},{\bm{A}}^{\prime},{\bm{\Omega}})$.
8:     if $|p-p_{0}|<\epsilon$ then
9:        ${\sigma_{q}}_{\text{min}}\leftarrow\sigma_{q}$.
10:     else
11:        ${\sigma_{q}}_{\text{max}}\leftarrow\sigma_{q}$.
12:     end if
13:  end while
13:  $\{{\bm{B}}^{\prime},{\bm{A}}^{\prime},{\bm{\Omega}}\}$: Bayesianized Low-Rank Adapter.
@@ 4 Theoretical Analysis
We present our theoretical analysis in this section (with complete proofs provided in Appendix A). First, we demonstrate that our Bayesianization scheme, defined in Equations 4, 5, and 6, projects a full-rank isotropic Gaussian distribution onto the low-rank space. We then prove that optimizing Eqn. 7 is equivalent to performing variational inference for LLM’s weights under specific, achievable conditions.
@@
The evaluation metric $l_{\mathcal{D}}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is the Negative Log Likelihood (NLL) evaluated on the data distribution ${\mathcal{D}}$ for the given posterior standard deviation $\sigma_{q}$:
@@
Furthermore, we assume $l_{\mathcal{D}}$ is locally convex, i.e., there exists $\epsilon_{0}>0$ such that $l_{\mathcal{D}}^{\prime\prime}(\sigma_{q})>0$, for all $\sigma_{q}\in[0,\epsilon_{0})$.
@@
The local convexity of the loss function is not an unrealistic assumption (Milne, 2019). For instance, a local minimum ${\bm{W}}_{0}$ of a twice-differentiable loss function $l$ will imply the local convexity around ${\bm{W}}_{0}$, which has been widely assumed in Laplace Approximation (Tierney & Kadane, 1986; Bishop, 2006).
Theorem 4.1.
With the pre-trained weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, the low-rank weight update matrix $\{{\bm{B}}^{\prime}\in\mathbb{R}^{m\times r},{\bm{A}}^{\prime}\in\mathbb{R}^{r\times n}\}$ transformed from the given matrices $\{{\bm{B}},{\bm{A}}\}$ following Eqn. 4 and 5, suppose that the variational distribution of ${\bm{A}}^{\prime}$ is Gaussian $q({\bm{A}}^{\prime}|{\bm{\theta}})=\prod_{ij}{\mathcal{N}}(A_{ij}|M_{ij},\Omega_{ij}^{2})$, where ${\bm{M}}=[M_{ij}=A^{\prime}_{ij}]\in\mathbb{R}^{r\times n}$ is its mean and ${\bm{\Omega}}=[\Omega_{ij}]\in\mathbb{R}^{r\times n}$ is the standard deviation calculated as in Eqn. 6. The equivalent variational distribution $q(\operatorname{vec}({\bm{W}})|\sigma_{q})$ defined on the full weight matrix is
@@
${\bm{\Sigma}}_{q}$ $=\sigma_{q}^{2}\cdot{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &{\bm{O}}_{m-r}\end{bmatrix}.$
Theorem 4.1 establishes that for any given $\sigma_{q}$, our algorithm for regrouping ${\bm{B}},{\bm{A}}$ and computing the standard deviation matrix ${\bm{\Omega}}$ successfully constrains the corresponding full-weight variational distributions to the family of low-rank isotropic Gaussian distributions. While alternative families of Gaussian distributions parameterized by a single scale $\sigma_{q}$ are possible, our empirical results demonstrate that our proposed approach achieves superior performance (Sec. 5.3).
Theorem 4.2.
Suppose the evaluation metric $l_{\mathcal{D}}(\sigma_{q})$ defined following Assumption 4.1 is locally convex within the range of $\sigma_{q}\in[0,\epsilon_{0})$. Suppose the posterior distribution of ${\bm{W}}$ given $\sigma_{q}$ is defined following Theorem 4.2. Suppose we have the prior distribution $P(\operatorname{vec}({\bm{W}}))={\mathcal{N}}(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{p},{\bm{\Sigma}}_{p})$, where ${\bm{\mu}}_{p}={\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}})$, and ${\bm{\Sigma}}_{p}=\sigma_{p}^{2}{\bm{I}}$ with $\sigma_{p}\gg\epsilon_{0}$. Then $\exists\epsilon>0$, s.t. the following two optimizations
(i) Minimization of Variational Free-Energy (Blundell et al., 2015)
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\epsilon,$
are equivalent, where $\lambda$ is the coefficient of the KL-divergence regularization.
This theorem provides the primary theoretical foundation for our approach. It demonstrates that under specific conditions—namely, local convexity within $[0,\epsilon_{0})$ and prior standard deviation $\sigma_{p}\gg\epsilon_{0}$—maximizing the scale $\sigma_{q}$ of the standard deviation matrix is equivalent to minimizing the variational free energy, which corresponds to performing variational inference for the weight posterior distribution.
@@
TFB maintains theoretical soundness (through its equivalence to Variational Inference (VI)) while offering practical simplicity, as it eliminates the need to explicitly specify the prior distribution’s standard deviation $\sigma_{p}$. The condition $\sigma_{p}\gg\epsilon_{0}$ is naturally satisfied by common choices such as the standard normal distribution ($\sigma_{p}=1$) or uniform distribution ($\sigma_{p}\rightarrow+\infty$).
@@ 5 Experiments
In this section, we evaluat TFB through comprehensive experiments on various settings. Sec. 5.1 details our experimental setup, including baselines and evaluation protocols. We then present our main results on in- and out-of-distribution performance in Sec. 5.2, followed by an ablation study analyzing key components of our method in Sec. 5.3. To demonstrate the broad applicability of our approach, we conduct additional experiments on different LLM architectures (Sec. 5.4) and explore its compatibility with other LoRA-like PEFT methods (Sec. 5.5).
Table 1: Performance of different methods applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. “TF?” denotes whether a method is Training-Free. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 gradient steps. We sample $N=10$ during inference in all sampling-based methods including BLoB (Wang et al., 2024d) and TFB. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. For TFB, the anchor dataset ${\mathcal{D}}$ is set to a randomly sampled subset of the original training set, the performance evaluation metric $l$ is set to accuracy, and the performance drop tolerance is set adaptively to 1% or 0.5% based on whether the given mean overfits. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.

### v2

@@ 3.3 TFB as Variance Maximization
The previous section presents a straightforward Bayesianization scheme for a predetermined value of $\sigma_{q}$. In this section, we describe a practical method for determining $\sigma_{q}$.
@@
A General Bayesianization Framework. Consider an in-distribution “anchor” dataset ${\mathcal{D}}$, an associated evaluation metric $l$, and a performance change tolerance $\epsilon$. TFB determines $\sigma_{q}$ by solving a constrained optimization problem:
@@
where $l({\mathcal{D}}|{\bm{B}},{\bm{A}})$ and $l({\mathcal{D}}|{\bm{B}}^{\prime},{\bm{M}},{\bm{\Omega}}(\sigma_{q}))=\mathbb{E}_{{\bm{E}}\sim{\mathcal{N}}(0,{\bm{\Omega}}^{2})}[l({\mathcal{D}}|{\bm{B}}^{\prime},{\bm{M}}+{\bm{E}})]$ denote the pre- and post-Bayesianization performance, respectively. This optimization maximizes the noise scale $\sigma_{q}$ applied to model weights ${\bm{M}}$ while ensuring that the resulting performance change remains within an acceptable threshold $\epsilon$.
@@
Anchor Dataset ${\mathcal{D}}$ and Evaluation Metric $l$. Our general TFB framework accommodates various choices of anchor dataset ${\mathcal{D}}$ and evaluation metric $l$ based on practical requirements. Below, we consider two key scenarios (with $N$ being slightly overloaded in its notation).
@@
For supervised dataset ${\mathcal{D}}=\{{\bm{x}}_{n},y_{n}\}_{n=1}^{N}$: The Negative-Log Likelihood (NLL) serves as a natural evaluation metric in Eqn. 8: $l_{\text{nll}}({\mathcal{D}}|{\bm{\theta}})=-\tfrac{1}{N}\sum_{n=1}^{N}\log P_{\bm{\theta}}(y_{n}|{\bm{x}}_{n})$, as it theoretically corresponds to minimizing the KL-regularized variational objective (more details in Sec. 4). The anchor dataset ${\mathcal{D}}$ can be either the original training set used for the LoRA model or an independent calibration dataset, as commonly employed in calibration-based methods [16, 69]. Alternative evaluation metrics such as accuracy or F1 score are also readily applicable. In our experimental setup, to ensure fair comparisons across uncertainty estimation baselines, we use the original training data as ${\mathcal{D}}$ (maintaining the same information access as baselines) and employ NLL as the evaluation metric. Additional results with accuracy as $l$ can be found in Appendix F.5.
@@
For unsupervised dataset ${\mathcal{D}}=\{{\bm{x}}_{n}\}_{n=1}^{N}$: One approach is to generate pseudo-labels $\widehat{y}$ using the model before Bayesianization, effectively converting the problem to the supervised case with ${\mathcal{D}}=\{{\bm{x}}_{n},\widehat{y}_{n}\}_{n=1}^{N}$. TFB can also directly incorporate purely unsupervised metrics such as the expected embedding norm $l_{\text{emb}}({\mathcal{D}}|{\bm{\theta}})=\mathbb{E}_{{\bm{x}}\sim{\mathcal{D}}}[\|\text{emb}({\bm{x}}|{\bm{\theta}})\|]$, where we are only concerned with properties of the representations themselves rather than any supervised signal. Hence our TFB offers substantially more flexibility compared to pure calibration methods, which typically rely on a labeled unseen calibration dataset. As a general framework, TFB also supports alternative evaluation metrics and statistical measures specifically designed for unsupervised data.
Performance Change Tolerance $\epsilon$. The selection of performance change tolerance $\epsilon$ is critical in TFB. While our experiments demonstrate that a fixed relative change rate, i.e., $\nicefrac{{\epsilon}}{{p_{0}}}=0.3\%$ for NLL and $\nicefrac{{\epsilon}}{{p_{0}}}=1\%$ for accuracy, where $p_{0}$ denotes the pre-Bayesianization performance, can achieve effective uncertainty estimation across various datasets and LoRA checkpoints, an adaptive $\epsilon$ can further improve the performance of TFB. Users can determine the appropriate value for $\epsilon$ by considering multiple factors simultaneously, among which the most important is the given LoRA checkpoint. For instance, an overfitted LoRA can typically accommodate a larger tolerance $\epsilon$ when using the training dataset (or its subset) as the anchor dataset. Additional properties of the data, model, and adaptation tasks can inform the choice of $\epsilon$ as well.
@@ 3.4 TFB: Final Algorithm
Final TFB Algorithm: Automatically Determining $\sigma_{q}$. Our final algorithm, presented in Algorithm A and Fig. 1, employs binary search to determine the optimal $\sigma_{q}^{*}$ within an initial range $[{\sigma_{q}}_{\text{min}},{\sigma_{q}}_{\text{max}}]$. After identifying the optimal $\sigma_{q}^{*}$, we Bayesianize all LoRA layers using this value.
Prediction. For prediction, we average multiple outputs produced by samples from TFB’s posterior:
$P_{{\bm{\theta}}}(y|{\bm{x}})$ $=\mathbb{E}_{q({\bm{W}}|{\bm{\theta}})}[P(y|{\bm{x}},{\bm{W}})]\approx\tfrac{1}{N}\sum\nolimits_{n=1}^{N}P({\bm{y}}|{\bm{x}},{\bm{W}}_{n}),\quad{\bm{W}}_{n}\sim q({\bm{W}}|{\bm{\theta}}),$ (9)
where $q({\bm{W}}|{\bm{\theta}})$ denotes the variational distribution defined in Eqn. 3, and we set the number of test-time samples to $N=10$, following BLoB’s protocol [61].
Remark on TFB’s Efficiency. While TFB with binary search is efficient in terms of both time and memory (Appendix F.2), and yields near-optimal solution of $\sigma_{q}$ [^3] While traditional search algorithms require monotonicity within the search range to guarantee optimal solutions, empirically a near-optimal $\sigma_{q}$ is sufficient for effective uncertainty estimation., more efficient parallel searching technique can be applied in practice. For instance, in Appendix F.10, we conduct a grid search across 8 different $\sigma_{q}$ values in parallel, construct an approximate function $\widehat{\sigma}_{q}(p)$ through piecewise linear interpolation of the observed performance, and estimate $\sigma_{q}^{*}\approx\widehat{\sigma}_{q}(p_{0}-\epsilon)$, where $p_{0}$ denotes the model’s performance before TFB.
@@ 4 Theoretical Analysis
In this section, we discuss our theoretical analysis, with complete proofs provided in Appendix D. First, we demonstrate that our TFB’s Bayesianization scheme, defined in Equations 4, 5, and 7, projects a full-rank isotropic Gaussian distribution onto the low-rank space. We then prove that Eqn. 8 is equivalent to generalized variational inference for LLMs’ weights under specific, achievable conditions, offering solid theoretical grounding for TFB.
@@
The evaluation metric $l_{\mathcal{D}}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is the Negative Log Likelihood (NLL) evaluated on the data distribution ${\mathcal{D}}$ for the variational standard deviation $\sigma_{q}$:
@@
Furthermore, we assume $l_{\mathcal{D}}$ is locally convex, i.e., $\exists\epsilon_{0}>0$ s.t. $l_{\mathcal{D}}^{\prime\prime}(\sigma_{q})>0$, $\forall\sigma_{q}\in[0,\epsilon_{0})$.
@@
The local convexity of the loss function is not unrealistic [43]. For instance, a local minimum ${\bm{W}}_{0}$ of a twice-differentiable loss function $l$ will imply the local convexity around ${\bm{W}}_{0}$, as assumed in Laplace Approximation [52, 6].
Theorem 4.1 (Equivalent Variational Distribution of the Full Weight ${\bm{W}}$ in TFB).
With the pre-trained weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, the low-rank weight update matrix $\{{\bm{B}}^{\prime}\in\mathbb{R}^{m\times r},{\bm{A}}^{\prime}\in\mathbb{R}^{r\times n}\}$ transformed from the given matrices $\{{\bm{B}},{\bm{A}}\}$ following Eqn. 4 and 5, suppose that the variational distribution of ${\bm{A}}^{\prime}$ is Gaussian $q({\bm{A}}^{\prime}|{\bm{\theta}})=\prod_{ij}{\mathcal{N}}(A_{ij}|M_{ij},\Omega_{ij}^{2})$, where ${\bm{M}}=[M_{ij}=A^{\prime}_{ij}]\in\mathbb{R}^{r\times n}$ is its mean and ${\bm{\Omega}}=[\Omega_{ij}]\in\mathbb{R}^{r\times n}$ is the standard deviation calculated as in Eqn. 7. The equivalent variational distribution $q(\operatorname{vec}({\bm{W}})|\sigma_{q})$ defined on the full weight ${\bm{W}}$ is
@@
${\bm{\Sigma}}_{q}$ $=\sigma_{q}^{2}{\bm{I}}_{n}\otimes\begin{bmatrix}{\bm{I}}_{r}&\\ &\mathbf{0}_{m-r}\end{bmatrix}.$
Theorem 4.1 establishes that for any given $\sigma_{q}$, our algorithm for regrouping ${\bm{B}},{\bm{A}}$ and computing the standard deviation matrix ${\bm{\Omega}}$ successfully constrains the corresponding full-weight variational distributions to the family of low-rank isotropic Gaussian distributions. This lays the foundation for the equivalence between our TFB and generalized variational inference to approximate the posterior distribution of LLM parameters (details in Theorem 4.2).
While alternative families of Gaussian distributions parameterized by a single scale $\sigma_{q}$ are possible, our empirical results demonstrate that our approach achieves superior performance (Sec. 5.3).
Theorem 4.2 (TFB as Generalized Variational Inference).
Suppose the evaluation metric $l_{\mathcal{D}}(\sigma_{q})$ defined following Assumption 4.1 is locally convex within the range of $\sigma_{q}\in[0,\epsilon_{0})$. Suppose the approximate distribution of ${\bm{W}}$ given $\sigma_{q}$ is defined following Theorem 4.1. Suppose we have the prior distribution $P(\operatorname{vec}({\bm{W}}))={\mathcal{N}}(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{p},{\bm{\Sigma}}_{p})$, where ${\bm{\mu}}_{p}={\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}^{\prime}{\bm{M}})$, and ${\bm{\Sigma}}_{p}=\sigma_{p}^{2}{\bm{I}}$ with $\sigma_{p}>\epsilon_{0}$. Then for $\forall\lambda>0$, $\exists\widetilde{\epsilon}$, s.t. the following two optimization problems (i) Generalized Variational Inference [7, 22, 30, 34]
@@
s.t. $l_{\mathcal{D}}(\sigma_{q})\leq\widetilde{\epsilon},$
are equivalent, i.e., the two optimization problems have the same optimal solution, where $\lambda$ is the regularization coefficient of the KL-divergence.
This theorem provides the primary theoretical foundation for TFB. It demonstrates that under specific conditions – namely, local convexity within $[0,\epsilon_{0})$ and prior standard deviation $\sigma_{p}>\epsilon_{0}$ – maximizing the scale $\sigma_{q}$ of the standard deviation matrix is equivalent to generalized variational inference [34], which approximates the posterior distribution of LLM parameters. Notably, when $\lambda=\nicefrac{{1}}{{|{\mathcal{D}}|}}$ is set to the reciprocal of the dataset size, generalized variational inference reduces to variational inference.
@@
TFB maintains theoretical soundness (through its equivalence to variational optimization) while offering practical simplicity, as it eliminates the need to explicitly specify the prior distribution’s standard deviation $\sigma_{p}$. The condition $\sigma_{p}>\epsilon_{0}$ is naturally satisfied by common choices such as the standard normal distribution ($\sigma_{p}=1$) or uniform distribution ($\sigma_{p}\rightarrow+\infty$).
@@ 5 Experiments
We evaluate TFB through comprehensive experiments.
5.1 Settings
Models, Datasets, and Evaluation. We use the latest open-source Meta-Llama-3.1-8B as our primary LLM backbone while also providing additional results on other recent LLM architectures in Sec. 5.4, including llama-2-7b-hf, Meta-Llama-3-8B, and Mistral-7B-v0.3 from the Llama [12] and Mistral [26] families.
Table 1: Performance of different methods applied to LoRA on Llama3.1-8B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. “TF?” denotes whether a method is Training-Free. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after fine-tuning of 5 epochs. We use $N=10$ samples during inference in all sampling-based methods including BLoB [61] and TFB. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. Cells highlighted in green indicate improved performance achieved by TFB compared to the weight mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.